OpenStudy (anonymous):

what is the answer in 64-x3y3

OpenStudy (anonymous):

You can write the above expression as: \[4^3 - (xy)^3\] Use: \[a^3 - b^3 = (a-b)(a^2 + b^2 + ab)\] Here, a = 4 and b = xy Just plug in the values for a and b..